# Neural Context Flows for Meta-Learning of Dynamical Systems

### v4

@@ Appendix B Experimental Details
In this section, we share crucial details that went into the experiments conducted in Section 4 and expanded upon in Appendix C. Unless otherwise specified, the hyperparameters in the next 3 paragraphs were applied to produce all figures and tables in this work.
@@
Each method was tuned to provide an excellent balance of accuracy and efficiency. The main motivation behind our choices was to restraint model sizes both to allow efficient training of all baselines, and fair comparison at roughly equal parameter counts. Irreconcilable differences with the baseline adaptation rules make it difficult to perform a systematic comparison, which is why we focused on parameter count as done in Qin et al. ; all the while noting the all models involved had sufficient capacity to approximate the problems at hand. For NCF and CAVIA, we count the number of learnable parameters in the main network being modulated. For CoDA, we count the root network plus the hypernetwork’s learnable parameters as outlined in Equation 3.
@@
Our main workstation for training and adaptation was fitted with an Nvidia GeForce RTX 4080 graphics card which was used for the SP, LV, SM, and NS problems. Additionally, we used an RTX 3090 GPU of the same generation for the GO problem, and an NVIDIA A100 Tensor Core GPU for the GS problem as its CNN-based architecture was more memory-intensive. In addition to the training hardware and deep learning frameworks (JAX for NCFs, and PyTorch for CAVIA and CoDA) that made apple-to-apple comparison challenging, our network architectures varied slightly across methods, as the subsections below will highlight. This is part of the reason we decided to focus on total parameter count as the great equalizer.
@@
All in all, to ensure fair comparison, we made sure each model was sufficiently large to represent the task at hand (by comparing to commonly used hyperparameters in the literature, e.g., in Kirchmeyer et al. [2022]). Since a full iteration means something different for the frameworks under comparison, we simply made sure the models were trained for sufficiently long, i.e. the lowest-performing method for each task was trained for at least as long as the second-best for that task, and its best checkpoint restored at the end.
B.1 NCFs
@@
The MLPs used as our context, state and main networks (as depicted in Figure 2b) typically have a uniform width. When that is not the case, we use the notation $[h_{\text{in}}\rightarrow h_{1}\rightarrow...\rightarrow h_{d-1}\rightarrow h_{\text{out}}]$ to summarize the number of units in each layer of an MLP of depth $d$ for brevity. We use a similar notation to summarize the number of channels in a Convolutional Neural Network (CNN). For instance, on the (SP) problem, the context network was the MLP $[256\rightarrow 64\rightarrow 64\rightarrow 64]$ (which indicates a context size of $d_{\xi}=256$), the state network was $[2\rightarrow 64\rightarrow 64\rightarrow 64]$; the main network $[128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$ ; all with Swish activations Ramachandran et al. [2017]. Bellow, we follow the same convention to detail the networks used for all problems in this paper’s main comparison (see Table 1).
LV (context) $[1024\rightarrow 256\rightarrow 256\rightarrow 64]$; (state) $[2\rightarrow 64\rightarrow 64\rightarrow 64]$; (main) $[128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GO (context) $[256\rightarrow 64\rightarrow 64\rightarrow 122]$; (state) $[7\rightarrow 122\rightarrow 122\rightarrow 122]$; (main) $[244\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 7]$SM (context, state, main) Exact same architectures as with SP above • BT (context) linear layer with 256 input and 128 outputs features followed by an appropriate reshaping then a 2D convolution with 8 channels and circular padding to maintain image size, a kernel of size $3\times 3$, and Swish activation [^6] Except for the number of channels, these are default convolution settings we use throughout. The circular padding enforces the periodic boundary condition observed in the trajectory data.; (state) appropriate reshaping before 2D convolution with 8 channels; (main) CNN $[16\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GS Same network architectures as for the previous BT problem, with the exception that the linear layer in the context network outputs 2048 features • NS (context) A single linear layer with 202 input and 1024 output features reshaped and concatenated to the state and the grid coordinates before processing; (state) No state network is used for this problem; (main) 2-dimensional Fourier Neural Operator (FNO) Li et al. [2020] with 4 spectral convolution layers, each with 8 frequency modes and hidden layers of width 10. Its lifting operator is a $1\times 1$ convolution with 4 input and 10 output channels respectively – functionally similar to a fully connected layer with weight sharing. The projection operator has two such layers, with 10 inputs, 1 output, and 16 hidden channels in-between.
@@
As for the propagation of gradients from the loss function to the learnable parameters in the right-hand-side of the neural ODEs (5), we decided to differentiate through our solvers, rather than numerically computing adjoints Chen et al. [2018]. Since our main solver were the less demanding RK4 and Dopri5 solvers, we didn’t incur the heavy memory cost Differentiable Programming methods are typically known for Nzoyem et al. [2023], Kidger et al. [2020].
We use the integration schemes: Dopri5 Wanner and Hairer [1996] for GO, SM, GS, and BT, with relative and absolute tolerances of $10^{-3}$ and $10^{-6}$ respectively; RK4 for LV with $\Delta t=0.1$; Explicit Euler for NS with $\Delta t=1$. Across all NCF variants, we kept the context pool size under 4 to reduce computational workload ($p=2$ for LV and SM, and $p=4$ for LV and GO, and $p=3$ for all PDE problems). We use the Adam optimizers Diederik [2014] for both model weights and contexts on ODE problems, and Adabelief Zhuang et al. [2020] for PDE problems. Their initial learning rates were as follows: $3\times 10^{-4}$ for LV and NS, $10^{-3}$ for GO, BT, and GS, $10^{-4}$ for SM. That learning rate is kept constant throughout the various trainings, except for BT, GS, and NS where it is multiplied by a factor (0.1, 0.5, and 0.1 respectively) after a third of the total number of training steps, and again by the same factor at two-thirds completion. The same initial learning rates are used during adaptation, with the number of iterations typically 1500.
As per Equation 10, the NCF implementation is batched across both environments and trajectories for distributed and faster training. This is in contrast to the two baselines, in which predictions are only batched across trajectories. We use full batches to accelerate training. This means that with LV for instance, all 4, 32, or 1 will be used at once depending on the data split we use (meta-trainingmeta-testing, supportquery). For regularization of the loss function Equation 10, we set $\lambda_{1}=10^{-3}$ for all problems; but $\lambda_{2}=0$ for ODE problems and $\lambda_{2}=10^{-3}$ for PDE problems. For NCF-$\mathit{t}_{2}$ we always use a proximal coefficient of $\beta=10$ except for the LV where $\beta=100$.
The simple pendulum SP problem used in this work leveraged NCF-$\mathit{t}_{2}$. The integrator was Dopri5 with default tolerances, a proximal coefficient $\beta=10^{2}$, a constant learning rate of $10^{-4}$, 500 outer and 20 inner iterations [^7] The inner iterations used in NCF-$\mathit{t}_{2}$ are not to be confused with the inner gradient update steps in CAVIA, since NCF performs alternating rather than bi-level optimization.. For the other problems highlighted in Table 1, the remaining crucial hyperparameters are defined in Tables 3 and 4.
@@
# Inner iterations 25 10 10 20 20 25
@@
B.2 CoDA
The reference CoDA implementation from Kirchmeyer et al. [2022] was readily usable for most problems. We adapted its data generation process to incorporate two new benchmarks introduced in this paper (the SM and BT problems).
@@ B.2 CoDA
Neural networks without physics priors are employed throughout, all with the Swish activation function. • LV: 4-layers MLP with 224 neurons per hidden layers (width) • GO: 4-layers MLP with width 146 • SM: 4-layers MLP with width 90 • BT: 4-layers ConvNet with 46 hidden convolutional filters and $3\times 3$ kernels. Its outputs is rescaled by $10^{-4}$ to stabilize rollouts • GS: Same ConvNet as BT, but with 106 filters. • NS: 2-dimensional Fourier Neural Operator (FNO) Li et al. [2020] with 4 spectral convolution layers, each with 8 frequency modes and hidden layers of width 10. Its lifting operator is a single fully connected layer, while its projection operator has two such layers, with 16 hidden neurons.
@@
Using TorchDiffEq Chen [2018], CoDA backpropagates gradients through the numerical integrator. We use the same integration schemes as NCF above. It stabilizes its training by applying exponential Scheduled Sampling Bengio et al. [2015] with constant $k=0.99$ and initial gain $\epsilon_{0}=0.99$ updated every 10 epochs (except for the NS problem updated every 15 epochs). The Adam optimizer with a constant learning rate of $10^{-4}$ is used. The batch size and the number of epochs is given in the Table 5.
@@
In line with CoDA’s fundamental low-rank assumption, we maintain a context size $d_{\xi}=2$ for all problems, since no more than 2 parameters vary for any given problem.
B.3 CAVIA
The reference implementation of CAVIA-Concat Zintgraf et al. [2019] required substantial modifications to fit dynamical systems. Importantly, we incorporated the TorchDiffEq Chen [2018] open-source package and we adjusted other hyperparameters accordingly to match NCF and CoDA on parameter count and other relevant aspects for fair comparison. The performance of CAVIA-Concat is heavily-dependent on the context size, which we adjust depending on the problem.
@@ B.3 CAVIA
Like CoDA, the Swish activation is used for all neural networks • LV: 4-layers MLP with width 278 • GO: 4-layers MLP with width 168 • SM: 4-layers MLP with width 84 • BT: 7-layers ConvNet with 46 hidden convolutional filters and $3\times 3$ kernels. Its outputs is rescaled by $0.1$ to stabilize rollouts • GS: 4-layers ConvNet with 184 hidden convolutional filters and $3\times 3$ kernels. Its outputs is rescaled by $0.1$ to stabilize rollouts. • NS: Same FNO as with NCF above, including the single-layer context network with 1024 output neurons.
@@
We use full batches to accelerate training. We use the Adam optimizer with a learning rate of 0.001 for the meta-update step. The single inner update step has a learning rate of 0.1. The number of iterations is given in Table 6.

### v5

@@ Appendix D Experimental Details
This section shares crucial details that went into the experiments conducted in Section 4 and expanded upon in Appendices B and C. Unless otherwise specified, the hyperparameters in the next 3 paragraphs were applied to produce all figures and tables in this work.
@@
Each method was tuned to provide an excellent balance of accuracy and efficiency. The main motivation behind our choices was to restrain model sizes both to allow efficient training of all baselines and fair comparison at roughly equal parameter counts. Irreconcilable differences with the baseline adaptation rules made it difficult to perform a systematic comparison, which is why we focused on parameter count as done in Qin et al. ; all the while noting that all models involved had sufficient capacity to approximate the problems at hand. For NCF and CAVIA, we counted the number of learnable parameters in the main network being modulated. For CoDA, we counted the root network plus the hypernetwork’s learnable parameters as outlined in Eq. 3.
@@
Our main workstation for training and adaptation was fitted with an Nvidia GeForce RTX 4080 graphics card which was used for the SP, LV, SM, and NS problems. Additionally, we used an RTX 3090 GPU of the same generation for the GO problem, and an NVIDIA A100 Tensor Core GPU for the GS problem as its CNN-based architecture was more memory-intensive. In addition to the training hardware and deep learning frameworks (JAX for NCFs, and PyTorch for CAVIA and CoDA) that made apple-to-apple comparison challenging, our network architectures varied slightly across methods, as the subsections below highlight. This is part of the reason we decided to focus on total parameter count as the great equalizer.
@@
To ensure fair comparison, we made sure each model was sufficiently large to represent the task at hand (by comparing to commonly used hyperparameters in the literature, e.g., in Kirchmeyer et al. (2022)). Since an “epoch” meant something different for the frameworks under comparison, we simply made sure the models were trained for sufficiently long, i.e., the lowest-performing method for each task was trained for at least as long as the second-best for that task, and its checkpoint with the best validation error was restored at the end.
D.1 NCFs
@@
The MLPs used as our context, state and main networks (as depicted in Fig. 2b) typically had a uniform width. When that was not the case, we use the notation $[h_{\text{in}}\rightarrow h_{1}\rightarrow...\rightarrow h_{d-1}\rightarrow h_{\text{out}}]$ to summarize the number of units in each layer of an MLP of depth $d$ for brevity. We used a similar notation to summarize the number of channels in a Convolutional Neural Network (CNN). For instance, on the SP problem, the context network was the MLP $[256\rightarrow 64\rightarrow 64\rightarrow 64]$ (which indicates a context size of $d_{\xi}=256$), the state network was $[2\rightarrow 64\rightarrow 64\rightarrow 64]$, and the main network was $[128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$, all with Swish activations Ramachandran et al. (2017). For the OFA and OPE experiments that do not require contextual information (cf. Tables 1 and 4), we delete the state and context networks then increase the hidden units of the main network to $156$ to match the NCF parameter count. Below, we follow the same convention to detail the networks used for other problems in this paper’s main comparison (see Table 2).
LV (context) $[1024\rightarrow 256\rightarrow 256\rightarrow 64]$; (state) $[2\rightarrow 64\rightarrow 64\rightarrow 64]$; (main) $[128\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GO (context) $[256\rightarrow 64\rightarrow 64\rightarrow 122]$; (state) $[7\rightarrow 122\rightarrow 122\rightarrow 122]$; (main) $[244\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 7]$SM (context, state, main) Exact same architectures as with SP above • BT (context) linear layer with 256 input and 128 output features followed by an appropriate reshaping then a 2D convolution with 8 channels and circular padding to maintain image size, a kernel of size $3\times 3$, and Swish activation [^12] Except for the number of channels, these were default convolution settings we used throughout. The circular padding enforced the periodic boundary condition observed in the trajectory data.; (state) appropriate reshaping before 2D convolution with 8 channels; (main) CNN $[16\rightarrow 64\rightarrow 64\rightarrow 64\rightarrow 2]$GS Same network architectures as for the previous BT problem, with the exception that the linear layer in the context network outputted 2048 features • NS (context) A single linear layer with 202 input and 1024 output features reshaped and concatenated to the state and the grid coordinates before processing; (state) No state network was used for this problem; (main) 2-dimensional Fourier Neural Operator (FNO) Li et al. (2020) with 4 spectral convolution layers, each with 8 frequency modes and hidden layers of width 10. Its lifting operator was a $1\times 1$ convolution with 4 input and 10 output channels respectively – functionally similar to a fully connected layer with weight sharing. The projection operator had two such layers, with 10 inputs, 1 output, and 16 hidden channels in-between.
@@
As for the propagation of gradients from the loss function to the learnable parameters in the right-hand-side of the neural ODEs (5), we opted to differentiate through our solvers, rather than numerically solving for adjoint states Chen et al. (2018). Since our main solvers were the less demanding RK4 and Dopri5 solvers, we didn’t incur the heavy memory cost Differentiable Programming methods are typically known for Nzoyem et al. (2023); Kidger et al. (2020).
Specifically, we used the following integration schemes: Dopri5 Wanner & Hairer (1996) for GO, SM, GS, and BT, with relative and absolute tolerances of $10^{-3}$ and $10^{-6}$ respectively; RK4 for LV with $\Delta t=0.1$; Explicit Euler for NS with $\Delta t=1$. Across all NCF variants, we kept the context pool size under 4 to reduce computational workload ($p=2$ for LV and SM, and $p=4$ for LV and GO, and $p=3$ for all PDE problems). We used the Adam optimizers Diederik (2014) for both model weights and contexts on ODE problems, and Adabelief Zhuang et al. (2020) for PDE problems. Their initial learning rates were as follows: $3\times 10^{-4}$ for LV and NS, $10^{-3}$ for GO, BT, and GS, $10^{-4}$ for SM. That learning rate was kept constant throughout the various trainings, except for BT, GS, and NS where it was multiplied by a factor (0.1, 0.5, and 0.1 respectively) after a third of the total number of training steps, and again by the same factor at two-thirds completion. The same initial learning rates were used during adaptation, with the number of iterations typically set to 1500.
As per Eq. 10, the NCF implementation was batched across both environments and trajectories for distributed and faster training. This was in contrast to the two baselines, in which predictions were only batched across trajectories. Using full batches to accelerate training meant that with LV for instance, all 4, 32, or 1 trajectories were used at once depending on the data split (meta-training vs. meta-testing, support vs. query). For regularization of the loss function Eq. 10, we set $\lambda_{1}=10^{-3}$ for all problems; but $\lambda_{2}=0$ for ODE problems and $\lambda_{2}=10^{-3}$ for PDE problems. For NCF-$\mathit{t}_{2}$ we always used a proximal coefficient $\beta=10$ except for the LV where $\beta=100$.
As for the simple pendulum SP problem used in this work, it leveraged NCF-$\mathit{t}_{1}$ with $p=4$ and $d_{\xi}=256$. The integrator was Dopri5 with default tolerances, a proximal coefficient $\beta=10^{2}$, a constant learning rate of $10^{-4}$, and 12000 epochs. The OFA and OPE comparisons used the same hyperparameters, but with 6000 epochs for OPE and 2000 for OFA. For the other problems highlighted in Table 2, the remaining crucial hyperparameters (not stated above) are defined in Tables 8 and 9.
@@
# Inner iterations [^13] The inner iterations used in NCF-$\mathit{t}_{2}$ are not to be confused with the inner gradient update steps in CAVIA, since NCF performs alternating rather than bi-level optimization. 25 10 10 20 20 25
@@
D.2 CoDA
The reference CoDA implementation from Kirchmeyer et al. (2022) was readily usable for most problems. We adapted its data generation process to incorporate two new benchmarks introduced in this paper (the SM and BT problems).
@@ D.2 CoDA
Neural networks without physics priors were employed throughout, all with the Swish activation function. • LV: 4-layers MLP with 224 neurons per hidden layers (width) • GO: 4-layers MLP with width 146 • SM: 4-layers MLP with width 90 • BT: 4-layers ConvNet with 46 hidden convolutional filters and $3\times 3$ kernels. Its output was rescaled by $10^{-4}$ to stabilize rollouts • GS: Same ConvNet as BT, but with 106 filters. • NS: 2-dimensional Fourier Neural Operator (FNO) Li et al. (2020) with 4 spectral convolution layers, each with 8 frequency modes and hidden layers of width 10. Its lifting operator was a single fully connected layer, while its projection operator had two such layers, with 16 hidden neurons.
@@
Using TorchDiffEq Chen (2018), CoDA backpropagates gradients through the numerical integrator. We used the same integration schemes as NCF above. We stabilized its training by applying exponential Scheduled Sampling Bengio et al. (2015) with constant $k=0.99$ and initial gain $\epsilon_{0}=0.99$ updated every 10 epochs (except for the NS problem updated every 15 epochs). The Adam optimizer with a constant learning rate of $10^{-4}$ was used. The batch size and the number of epochs are given in the Table 10.
@@
We note that in line with CoDA’s fundamental low-rank assumption, we maintained a context size $d_{\xi}=2$ throughout this work since no more than 2 parameters varied for any given problem.
D.3 CAVIA
The reference implementation of CAVIA-Concat Zintgraf et al. (2019) required substantial modifications to fit dynamical systems. Importantly, we incorporated the TorchDiffEq Chen (2018) open-source package and we adjusted other hyperparameters accordingly to match NCF and CoDA on parameter count and other relevant aspects for fair comparison. The performance of CAVIA-Concat was heavily dependent on the context size, which we adjusted depending on the problem.
@@ D.3 CAVIA
Like CoDA, the Swish activation was used for all neural networks • LV: 4-layers MLP with width 278 • GO: 4-layers MLP with width 168 • SM: 4-layers MLP with width 84 • BT: 7-layers ConvNet with 46 hidden convolutional filters and $3\times 3$ kernels. Its outputs was rescaled by $0.1$ to stabilize rollouts • GS: 4-layers ConvNet with 184 hidden convolutional filters and $3\times 3$ kernels. Its outputs was rescaled by $0.1$ to stabilize rollouts. • NS: Same FNO as with NCF above, including the single-layer context network with 1024 output neurons.
@@
We used full batches to accelerate training. We use the Adam optimizer with a learning rate of 0.001 for the meta-update step. The single inner update step had a learning rate of 0.1. The number of iterations is given in Table 11.